OpenStudy (anonymous):

Solve the following system of equations using the substitution method. Enter your answers as (x,y) pairs, separated by commas. x+3y=14 xy=11

OpenStudy (anonymous):

x= 14-3y (14-3y)y = 11 14y-3y^2 - 11 =0 3y^2-14y+11=0 3.6666666 , 1 x1 = 14-3*3.66666 = 3 x2 = 14 - 3 = 11 (3,3.6666) , (11,1) notice that 3.66666 - is 3 and 2 thirds

OpenStudy (anonymous):

I would solve the top equation for x x = 14-3y then substitute into the bottom (14-3y)y=11 14y-3y^2=11 Get all the terms together -3y^2+14y-11=0 Use quadratic formula to find y (you will get two values) Use those values to find the x values

hero (hero):

I already posted the solution for this. This question has been posted twice.

OpenStudy (anonymous):

That seems to be happening tonight a bunch

hero (hero):

1. Solve both equations for y: y = (14-x)/3 y = 11/x Set y = y: (14-x)/3 = 11/x Cross Multiply: x(14-x) = 3*11 14x - x^2 = 33 Put in Standard form: x^2 -14x + 33 = 0 Solve for x: x^2 - 11x - 3x + 33 = 0 x(x-11) - 3(x-11) = 0 (x-11)(x-3) = 0 x = 3, 11 y = 1, 11/3 (11,1) (3,11/3)

OpenStudy (anonymous):

Thank you all